The place value of a digit at the tenths place is 10 times the same digit at the ones place.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False.

Let us assume a digit be ‘y’.
The place value of a digit at the tenths place = y × (1/10)

= y/10

Then,
The tenths place is 10 times the same digit at the ones place. y/10 = 10y is not possible


Thanks